You have a portfolio worth $101,000 that has an expected return of 11.8 percent. The portfolio has $18,400 invested in Stock O, $26,200 invested in Stock P, with the remainder in Stock Q. The expected return on Stock O is 16.1 percent and the expected return on Stock P is 13 percent. What is the expected return on Stock Q?
Answer #1

Ans 9.84%

RETURN ON O + RETURN ON P + RETURN ON Q = TOTAL RETURN
18400 * 16.1% + 26200 * 13% + 56400 * r% = 101000 * 11.8%
6368.4 + 56400 * r% = 11918
r% = (11918 - 6368.4) / 56400
r% = 9.84%
INVESTMENT IN Q = TOTAL INVESTMENT- INVESTMENT IN O - INVESTMENT IN P
101000 - 18400 - 26200
56400